OpenStudy (anonymous):

An open top shipping crate with a square bottom and rectangular sides is to hold 32 inches cubed and requires a minimum amount of cardboard. Find the most economical dimensions.

OpenStudy (anonymous):

we can do this if you like, it is not too hard

OpenStudy (anonymous):

yeah it would be great if you helped me

OpenStudy (anonymous):

ok put the sides of the base as x, so the area of the base is \[x^2\] and the hight as h so the volume is \[x^2h\]

OpenStudy (anonymous):

and since you know the volume is 32, you know \[x^2h=32\] so \[h=\frac{32}{x^2}\] so now we can write the surface area only in terms of x

OpenStudy (anonymous):

now the surface area is 4 sides and the bottom. the bottom has area \[x^2\] and the sides each have area \[xh\] and there are four of them so the total area is \[x^2+4xh\] then replace h by \[\frac{32}{x^2}\] and get the surface area S as a function of x alone

OpenStudy (anonymous):

okay i got it from here. thank you

OpenStudy (anonymous):

i get \[S(x)=x^2+4\times \frac{32}{x^2}=x^2+\frac{128}{x}\]

OpenStudy (anonymous):

do you think you could help me through 2 more problems?

OpenStudy (anonymous):

check my algebra, but if it is right now you job is to minimize the surface area. take the derivative, set it equal to zero, solve, etc.

OpenStudy (anonymous):

sure you can ask here or post afresh either way.

OpenStudy (anonymous):

damn typo in this line it should have been \[S(x)=x^2+4x\times \frac{32}{x^2}=x^2+\frac{128}{x}\]